OpenStudy (anonymous):

90n+23p=4523 If n and p are positive integers in the equation above, what is one possible value of n+p?

OpenStudy (anonymous):

In this problem you would just plug in some of the possible answers.

OpenStudy (anonymous):

I would start with the answers that have small factors

OpenStudy (anonymous):

it's not multiple choice though, so is there any time efficient ways?

OpenStudy (anonymous):

*are there

OpenStudy (anonymous):

90n = 4523 -23p 90n = 4500 +23 -23p 90n = 4500 +23 (1-p) Now in this equation we can plug p = 1 and get n = 50 so one solution of n+p = 51

OpenStudy (anonymous):

We could work it further for more solutions 90n = 4500 +23(1-p) 90n-4500 = 23 (1-p) 90(n-50) = 23 (1-p) \[n-50 = \frac{23 (1-p)}{90}\] Since n and p are both integers n-50 should also be an integer Thus we may say that 23(1-p) should be a multiple of 90 this condition is met only when (1-p) = 0 or 90 or 180 or 270 and so on. so you may get as many values from this condition as you want.

OpenStudy (anonymous):

hmm ok cool thanks!